OpenStudy (anonymous):

What is the volume of a cereal box that is 2 inches thick, 12 inches high and 10 inches wide?

OpenStudy (er.mohd.amir):

simply volume=l*b*h=2*12*10=240 in^3

OpenStudy (anonymous):

240 in ^2

OpenStudy (kash_thesmartguy):

Easy: Volume of any rectangular figure = l*b*h Therefore: \[2\times12\times10=240^{3}inches\]

OpenStudy (study_together):

The volume is L*W*H and so it will be 12*2*10 and so the answer will be

OpenStudy (deving7240):

240 in cubed